Find the differential of the function. z = 7t- - 1 / 1 7 +2 dz = Submit Answer
Answer #1

Consider Z= H- X? Since dz = dz.dt at Now dz d dt (1+-) 7+ 14 da 73 So da= (1+1 13 ) et